OpenStudy (anonymous):

A box has dimensions of 17 inches long, 1.3 feet wide, and 8 inches high. What is the volume of the box? The formula for the volume is V = l • w • h. 2121.6 cubic inches (in3) 176.8 cubic inches (in3) 14.7 cubic inches (in3) 1.2 cubic inches (in3)

OpenStudy (anonymous):

whats the answer! i need help

OpenStudy (anonymous):

V = (17)(1.3)(8) V = 176.8

OpenStudy (anonymous):

wait...that might not be right...hold on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is the width in the problem supposed to be in feet or inches ?

OpenStudy (anonymous):

12/1 = x/1.3 cross multiply x = (12)(1.3) x = 15.6 V = L x W x H V = (17)(15.6)(8) V = 2121.6 I had to convert 1.3 ft to inches.